46. What is the percent ionization of a 0.491 M aqueous solution of carbonic acid? Ka1= 4.2x10^-7; Ka2= 4.8x10^-11

47. What is the percent ionization of a 0.356 M aqueous solution of hydrosulfuric acid? Ka1= 9.5x10^-8; Ka2= 1x10^-19

48. What is the pH of a 0.387 M aqueous solution of acetic acid? Ka (CH3COOH) = 1.8x10^-5

49. What is the pH of a 0.0398 M aqueous solution of benzoic acid? Ka (C6H5COOH) = 6.5x10^-5

50. What is the pH of a 0.415 M aqueous solution of hydrofluoric acid? Ka (HF) = 7.1x10^-4


Please Help & Show Work! Thanks!
Answer #1

(46) Formula to find percent ionization is : Percent ionization = [17] X 100 - THAIgnitial (C) change -n H2CO3 ( carbonic aci, and (HCO3] = x .. (16934.53110m At equilibrium (H+) = x 2 :[ht] = 4.53X1ÖLMt using - HCO H+ + cor- Initial 4.537104 4.538100.491M r Initial conce [HA] = Put these values in equo Percent impatiens 4053xjoy Percentionnation = -X100 = 0.092 % ✓ 0.497is 9.50.356x168 x = 338 xtö8 ar r= 13:38 xjö8 = 1.84x164 At equilibrium (H+) = (45) = i [H +] = [H Ś] = 1.89 X OM ✓ being @Sritied Conce, [HA] = 0.356 ; Percent ionigetim = 1.348109 xloo 01356 = 0.052016 Percent ionizatime 0.0520 | 43 Chz cool = Chphu give by При = -lba Cot] Bu =-lag (2163Xčo?) [iky axbelog at lg 6] PH=-{log 2063 + log [03] [log ab = bloga] pu= -(0.419 -PH=-log Cut] s-log (1.61% 103] =-(loytet thoros] =-[0.207 -3 ] = 2.79 [P4 = 2.79 | V [long log1= 0.207] 50 Initial change Equ